NEAR-INFRARED DOUBLE RESONANCE SPECTRUM OF THE CH_{2} b̃←ã ORIGIN BAND

ZHONG WANG, Department of Chemistry, The Stony Brook University, Stony Brook, New York 11794; BOR-CHEN CHANG, Department of Chemistry, National Central University, Jhongli 32001,Taiwan; \underline{TREVOR J. SEARS}, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973 and Department of Chemistry, The Stony Brook University, Stony Brook, New York 11794.

Since Herzberg and Johns reported the first observation of the CH_{2} b̃←ã singlet-singlet transition, there have been numerous experimental and theorectical investigations of this system. Although the spectrum of the CH_{2} b̃←ã transition has now been studied from the ultraviolet to the near-infrared (NIR), the origin band has not previously been observed due to a poor Franck-Condon factor as well as the lack of a convenient high resolution light source at approximately 1.2 μm. Here we report the successful observation of the origin band of the CH_{2} b̃←ã system using a NIR-NIR double resonance technique. Based upon our results, effective rotational constants for the b̃ (0,0,0) level can be determined. The level positions also provide information that can be used to improve the potential energy surface for the b̃ state. Singlet CH_{2} is one of very few polyatomic molecules for which continuous high resolution data exist for rovibronic levels from the zero-point level of the ã state through to predissociation. Acknowledgments: This work at Brookhaven National Laboratory was carried out under Contract No.DE-AC02-98CH10886 with the U.S. Department of Energy, and BCC is supported by National Science Council, Taiwan.